OpenStudy (anonymous):

Find dy/dx given 13√x + 8√y = 3. Differentiate both sides of the equation with respect to x, and then solve for dy/dx. Do not substitute for y after solving for dy/dx. dy/dx=_______ Super confused:( thank you!!

OpenStudy (anonymous):

you're doing an implicit differentiation?

zepdrix (zepdrix):

Soooo what are you stuck on? :U The y term giving you trouble?

OpenStudy (anonymous):

@celestialdictator i believe so... not sure the exact term and @zepdrix i'm just not too sure where to start... am i setting it equal to zero?

OpenStudy (anonymous):

write your radicals as exponents

OpenStudy (anonymous):

well the derivative of a constant is always zero

OpenStudy (anonymous):

so that right hand side is zero right away

OpenStudy (anonymous):

ohh okay so would it be x^(1/13) and y^(1/8) ?

OpenStudy (anonymous):

\[13x^{\frac{ 1 }{ 2 }}+8y^{\frac{ 1 }{ 2 }}=0\]

OpenStudy (anonymous):

use the power rule to differentiate the left hand side but when you differentiate the y, you have to write an extra y

OpenStudy (anonymous):

ohh okay so from here, i rearrange it to look like this? \[\frac{ 13 }{ 2\sqrt{x} } + \frac{ 8 }{ 2\sqrt{y} } \frac{ dy }{ dx } = 0\]

OpenStudy (anonymous):

did i do that correctly?

zepdrix (zepdrix):

oo good job! :)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

then now, ask your self what is it that you wanted solved?

OpenStudy (anonymous):

hehe thanks :) so umm from here, would i isolate the dy/dx ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it's better to keep the radicals as exponents until the very end

OpenStudy (anonymous):

okay, so would it look like this? \[\frac{ dy }{ dx }= - \frac{ 13\times2\sqrt{y} }{ 8\times2\sqrt{x} }\] :/

zepdrix (zepdrix):

Looks good!! Simplifyyyy \c:/

OpenStudy (anonymous):

do i get \[-\frac{ 13\sqrt{y} }{ 8\sqrt{x} }\] ?

OpenStudy (anonymous):

you tell us :P

OpenStudy (anonymous):

haha umm i got that because the two 2's cancel out? would that be the final answer though? :/

zepdrix (zepdrix):

Yah looks fine. You can combine the roots if you want,\[\Large\bf\sf -\frac{13}{8}\sqrt{\frac{y}{x}}\]Looks a little nicer.

OpenStudy (anonymous):

looks a lot neater hehehe

OpenStudy (anonymous):

ahh okay awesome!!! thank you both!!! :)

OpenStudy (anonymous):

you didn't have to simplify, because it does not add or give any further information. it just makes equation look neater

OpenStudy (anonymous):

when I become a professor, I wouldn't even ask my students to simplify. Just give me the correct answer and I would be satisfied

OpenStudy (anonymous):

good job iheart and it's good that you're on calculus1 now

OpenStudy (anonymous):

ohh okay cool :) and yes!! i can assure you that the students will appreciate that!! :P thank you :)